# Divergence-Based Similarity Function for Multi-View Contrastive Learning

### v1

@@ Abstract
Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of an instance. While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views. In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions. Extensive experiments demonstrate that DSF consistently improves performance across various tasks, including kNN classification and linear evaluation, while also offering greater efficiency compared to other multi-view methods. Furthermore, we establish a theoretical connection between DSF and cosine similarity, and show that, unlike cosine similarity, DSF operates effectively without requiring a temperature hyperparameter.
@@ 1 Introduction
Figure 1: Illustration of feature‐distribution modeling on the unit hypersphere: given multiple augmented views of the image, their extracted feature vectors (dots) lie on the hyperspherical manifold and are used to construct smooth probability distributions (colored regions) over the sphere.
Learning effective representations without costly labels is a central challenge in modern machine learning. Self-supervised learning (SSL) addresses this by deriving supervisory signals from the data itself, and contrastive learning has emerged as one of its most successful paradigms. In contrastive frameworks such as SimCLR[Chen et al., 2020a], MoCo[He et al., 2020], and DCL[Yeh et al., 2022], representations of two augmented views of the same instance are ”pulled together” while being ”pushed apart” from all other samples. A key ingredient across these methods is cosine similarity on the unit hypersphere, which normalizes features to lie on a sphere and then measures their dot-product similarity.
Figure 2: Illustration of multi-view contrastive learning strategies. A data instance is augmented $M$ times, resulting in $x^{1},\cdots,x^{M}$. Base encoder outputs $M$ representations $z^{1},\cdots,z^{M}$ from $M$ views. From top to bottom: (i) Loss_avg, which averages pairwise loss terms across $M$ views; (ii) Fea_avg, which computes a mean embedding over the $M$ views before applying contrast; and (iii) DSF, our proposed model that aggregates the view’s feature into distribution and fuses these distributions into a single contrastive loss
However, most existing contrastive methods focus on pairwise relationships between two views. Recent efforts to leverage more than two views can be broadly categorized into two approaches.:
  • Loss-level approaches: methods that formulate the contrastive objective by averaging pairwise losses computed between all available combinations of views, typically using arithmetic or geometric averaging schemes, as in [Shidani et al., 2024].
  • Feature-level approaches: methods that first aggregate multi-view representations—most notably through feature averaging—before applying a contrastive objective. For example, FastSiam [Pototzky et al., 2022] computes the average of view features, while other techniques like M3G[Piran et al., 2024] adopt more complex strategies such as multi-marginal transport.
@@
To overcome these limitations, we introduce a novel similarity function that directly models the joint structure among all $M$ augmented views of an instance. Rather than reducing multi-view interactions to pairwise comparisons or expensive transport computations, our method first estimates a probability distribution over the views on the unit hypersphere and then computes the similarity between two instances via a negative divergence measure between their respective distributions. In particular, we employ the von Mises–Fisher distribution, which is parameterized by a mean direction $\mu$ and a concentration parameter $\kappa$. The mean direction $\mu$ represents the central tendency of the distribution, while the concentration parameter $\kappa$ captures the joint structure among the multi-view features. Moreover, under the special case of two views, our divergence based similarity exactly recovers the cosine similarity used in the InfoNCE loss, thereby bridging the gap between conventional pairwise and multi-view contrastive paradigms.
@@
  • Relation with original InfoNCE loss. We show that, when restricted to two views under specific condition, our similarity recovers cosine similarity in InfoNCE loss, thus offering a unified view of pairwise and multi-view contrastive methods.
  • Hyperparameter-free temperature control We mathematically establish the role of the temperature term in contrastive loss and show that our model learns optimally without manual temperature adjustment.
  • Experimental Results. We validate the effectiveness of our method through extensive experiments on kNN classification, linear evaluation, and resource analysis in terms of GPU memory and training time, consistently outperforming the state-of-the-art contrastive learning approaches.
The remainder of this paper is structured as follows. Section 2 provides an overview of contrastive learning and its multi-view extensions. Section 3 reviews related work on contrastive frameworks with both two-view ($M=2$) and multi-view ($M\geq 3$) settings. In Section 4, we present our proposed divergence-based similarity formulation. Section 5 describes the experimental setup and results. Finally, Section 6 concludes the paper and discusses potential directions for future research.
2 Preliminaries
@@
Figure 3: More Time Cost Comparison of performance over training time for kNN and linear evaluation. Both figures report accuracy for different training epochs and wall-clock time under a fixed GPU memory budget, with the number of views set to 8 ($M=2m=8$). The right-hand side of each figure shows the elapsed time of MoCo baselines, while the x-axis indicates the time and epochs for the multi-view methods. Note that the lowest dashed line (800 epochs, 5.5h) corresponds to the same GPU memory and training time as the first point(250 epochs, 5.5h) of each multi-view method. Compared to Loss_avg and Feature_avg at 800 epochs, our proposed DSF achieves more than 2$\times$ faster convergence in the kNN setting and over 3$\times$ faster convergence in the linear evaluation setting.
@@ 2.1 Contrastive Learning
Contrastive Learning (CL) has emerged as a powerful paradigm in unsupervised and self-supervised learning, aiming to learn representations by distinguishing positive and negative pairs. Here, positives refer to semantically similar pairs whereas negatives to dissimilar ones. The core idea is to map positive samples closer in the embedding space while pushing negative samples apart. Given a positive pair and a set of negative samples, this loss encourages the model to assign high similarity to the positive pair relative to the negatives. The loss can be formulated as follows:
$L_{i}=-\log\frac{e^{(sim(z_{i},z^{+}_{i})/\tau)}}{e^{(sim(z_{i},z^{+}_{i})/\tau)}+\sum\limits_{i\neq j}^{K}e^{(sim(z_{i},z^{-}_{j})/\tau)}}$ (1)
where $z^{+}_{i}$ is the embeddings of the positive pair respect to $z_{i}$ and $z^{-}_{j}$s are embeddings of $K-1$ negative pairs. $sim(\cdot,\cdot$) is a similarity measure (e.g., dot product) and $\tau$ is a temperature parameter.
@@
The von Mises-Fisher (vMF) distribution is a probability distribution dealing with a data that lies on a unit hypersphere. Defined over a $(p-1)$-dimensional unit hypersphere $\mathbb{S}_{p-1}$, vMF distribution describes the probability density of a unit vector with a specified mean direction. The distribution is characterized by a mean direction, indicating the central orientation of the vector, and a concentration parameter, which describes how tightly the distribution clusters around this mean direction.
The probability density function of the vMF distribution for a unit vector $x\in S_{p-1}$ is given by:
$f_{p}(x;,\mu,\kappa)=C_{p}(\kappa)\exp(\kappa\mu^{T}x)$ (2)
where $\mu$ is the mean direction, $\kappa$ is the concentration parameter, and $C_{p}(\kappa)$ is a normalization constant.
A higher concentration parameter $\kappa$ results in samples that are tightly clustered around the mean direction $\mu$, while $\kappa=0$ corresponds to a uniform distribution over the hypersphere.
3 Related Works
3.1 Original Contrastive Learning
Contrastive Learning appears in many self-supervised learning methods. SimCLR[Chen et al., 2020a] leverages the InfoNCE objective[Oord et al., 2018] to pull together representations of two augmented views while pushing apart all other samples. MoCo[He et al., 2020] maintains a momentum encoder, which is a moving average of the parameters to supply a large, consistent set of negatives without labels. [Yeh et al., 2022] decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable, high performance training with smaller batches and without a momentum encoder. MoCo v3 [Chen et al., 2021] explores contrastive learning using a transformer-based model, specifically the Vision Transformer (ViT) [Dosovitskiy et al., 2020].
3.2 Contrastive Learning with $M\geq 3$ views
Existing $M\geq 3$ multi‐view contrastive learning methods have broadly fallen into two categories. The first category shapes the loss by averaging pairwise terms across views. [Shidani et al., 2024] generate $M$ views per sample and aggregate their pairwise losses either arithmetically or geometrically. The second category is a method that deal with view representations at the feature level before computing the final objective. FastSiam[Pototzky et al., 2022] averages the other $M-1$ view features to define each view’s target , and M3G[Piran et al., 2024] adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost. In contrast to these approaches, our method integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
@@ 4 Divergence-based similarity function
In this section, we explain the details of proposed divergence-based similarity function(DSF) and relations with cosine similarity function. In Sec4.1, we formalize notations throughtout this paper. In Sec4.2, we describe how to estimate vMF distribution parameters $\mu,\kappa$. In Sec4.3,we propose the adjustment to kappa value for computational stabilization. In Sec4.4, we describe the calculation of DSF. In Sec4.5, we show that KL divergence-based similarity function between vMF distribution becomes equivalent to consine similarity function in InfoNCE Loss under certain conditions. Furthermore, we show why our divergence similairty function does not need a temperature hyperparameter and why cosine similarity function need temperature hyperparamter.
4.1 Notation
We denote the batch size, number of views, and feature dimensionality by $B$, $M$, and $D$, respectively. Each encoder is given $m$ views, where $m$ indicates the number of augmented views per encoder, resulting in a total of $M=2m$ views.
The Kullback–Leibler (KL) divergence is written as $D_{\mathrm{KL}}(\cdot\|\cdot)$, and we denote a von Mises–Fisher (vMF) distribution by the calligraphic symbol $\mathcal{D}$.
Let $x$ be an image instance, and $x^{l}$ denote its $l$-th augmented view. Given an encoder $f_{\theta}(\cdot)$, the corresponding feature on the unit hypersphere is computed as $z^{l}=f_{\theta}(x^{l})/\|f_{\theta}(x^{l})\|$, $l=1,\cdots,m$.
contrastive Learning(CL) makes $\mathbb{R}^{B\times 2\times D}$ features with two augmentation with each instance. Multi-view contrastive learning makes $\mathbb{R}^{B\times 2m\times D}$ features with total $M=2m$ views with each instance.
In CL, there are $2B$ features $z_{i}$, $i=1,\cdots,2B$. $z_{i}$ and $z_{B+i}$ are positively correlated. In Multi-view CL, there are $B\cdot 2m$ features $z_{i}^{l}$, $i=1,\cdots,B$ and $l=1,\cdots,m$. $\{z_{i}^{l}\}_{l=1}^{m}$ and $\{z_{B+i}^{l}\}_{l=1}^{m}$ are positively correlated.
Dataset Base Model Method Epoch kNN↑ Linear Evaluation↑
ImageNet-100 MoCo v3 - 300 74.08 81.52
MoCo v3 loss_avg 100 76.62 84.24
MoCo v3 fea_avg 100 76.94 84.06
MoCo v3 DSF 100 79.00 85.66
CIFAR10 MoCo - 800 88.44 88.10
MoCo loss_avg 250 88.28 88.58
MoCo fea_avg 250 87.47 88.02
MoCo DSF 250 90.04 91.21
Table 1: Comparison of kNN and Linear Evaluation Results Under Equivalent Resource Settings. All experiments are conducted under the same GPU memory and time constraints. The number of views is fixed at $M=2m=8$.
4.2 Estimate vMF Distribution
Before introducing the DSF, we first describe the process of constructing von Mises–Fisher (vMF) distributions from multiple augmented views of a single image instance.
Given an image instance $x$, we generate $M$ augmented views via data augmentation. These views are divided into two groups of $m$ samples each. All augmented images are passed through the encoder to obtain $M$ feature vectors on the unit hypersphere. Each group of features is then used to estimate the parameters of a vMF distribution. As a result, each instance is represented by two vMF distributions, each constructed from one group of augmented features.
Let $z^{1},\dots,z^{m}$ denote the feature vectors corresponding to the $m$ augmented views of an instance $x$. The mean direction $\mu$ and concentration parameter $\kappa$ of the vMF distribution are estimated as follows:
@@ 4.2 Estimate vMF Distribution
Since computing $\kappa=A_{p}^{-1}(\bar{R})$ is intractable in practice, we instead adopt an approximate form of $\kappa$ proposed by [Banerjee et al., 2005]:
$\kappa=\frac{\bar{R}(p-\bar{R}^{2})}{1-\bar{R}^{2}}$ (5)
Consequently, the vMF distribution constructed from the features $z^{1},\dots,z^{m}$ is denoted as:
$\mathcal{D}(z^{1},\dots,z^{m})=\mathcal{D}(\mu,\kappa).$ (6)
For subscript indexing $i$ :
$\mathcal{D}(z_{i}^{1},\dots,z_{i}^{m})=\mathcal{D}(\mu_{i},\kappa_{i})=\mathcal{D}_{i}.$ (7)
@@ 4.3 Stabilization of the Concentration Parameter $\kappa$
As a final step before introducing the DSF, we describe a strategy to stabilize the concentration parameter $\kappa$, which is essential for maintaining numerical stability during training.
As training progresses, the value of $\kappa$ can grow excessively, often leading to numerical overflow in computations. This issue arises from two primary factors:
  • High feature dimensionality. For a fixed mean resultant length $\bar{R}$, a larger feature dimension $p$ results in a sharper increase in $\kappa$. This becomes problematic, particularly when using higher-dimensional feature spaces (e.g., $D=128$ in our setup), as commonly done in larger models.
  • Feature concentration. As the encoder learns more discriminative representations, the extracted features become increasingly concentrated, causing $\bar{R}$ to approach 1, which in turn further amplifies $\kappa$.
Although this behavior is theoretically valid, it leads to numerical instability when computing divergence measures in practice. To mitigate this issue, we apply two simple yet effective techniques:
@@
    Dimension-based normalization: We normalize $\kappa$ by the feature dimension $p$, as $\kappa$ grows approximately linearly with $p$ when the mean resultant length $\bar{R}$ is fixed, as shown in Eq. 5. This prevents dimension-driven growth.
@@
    Scaling of mean resultant length: To avoid $\bar{R}$ approaching 1, we apply a scaling factor $\lambda_{\bar{R}}\in(0,1)$. In our experiments, we use $\lambda_{\bar{R}}=0.95$, modifying $\bar{R}$ as follows:
@@
These modifications are empirically shown to stabilize training and avoid divergence explosion in high-dimensional settings. Further details and illustrative figures can be found in the Appendix C Figure 6.

### v2

@@ Abstract
Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of data. While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views. In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions. Extensive experiments demonstrate that DSF consistently improves performance across diverse tasks—including kNN classification, linear evaluation, transfer learning, and distribution shift—while also achieving greater efficiency than other multi-view methods. Furthermore, we establish a connection between DSF and cosine similarity, and demonstrate that, unlike cosine similarity, DSF operates effectively without the need for tuning a temperature hyperparameter.
@@ 1 Introduction
Self-supervised learning (SSL) aims to learn meaningful representations without labels, often using contrastive learning (CL) frameworks such as SimCLR (Chen et al. 2020a), MoCo (He et al. 2020), and DCL (Yeh et al. 2022). These methods encourage similarity between augmented views of the same data while separating different data, typically via cosine similarity on unit-normalized features.
However, most existing contrastive methods focus on pairwise relationships between two views. Recent efforts to leverage more than two views can be broadly categorized into two approaches:
  • Loss-level approaches: Averaging pairwise losses computed between all available combinations of views (Shidani et al. 2024).
  • Feature-level approaches: Aggregating features (e.g., via averaging (Pototzky, Sultan, and Schmidt-Thieme 2022) or optimal transport (Piran et al. 2024)) before applying the contrastive loss.
Figure 1: Illustration of divergence-based similarity function: given multiple augmented views of the image, their extracted feature vectors (colored dots) lie on the hypersphere and are used to construct probability distributions (colored regions) over the unit hypersphere.
Figure 2: Illustration of multi-view contrastive learning strategies. Given an input $x$, we generate $M$ augmented views, resulting in $x^{1},\cdots,x^{M}$. Base encoder outputs $M$ representations $z^{1},\cdots,z^{M}$ from $M$ views. From top to bottom: (i) Loss_avg, which averages pairwise loss terms across $M$ views; (ii) Fea_avg, which computes a mean embedding over the $M$ views; and (iii) DSF, our proposed method, which aggregates features from $M$ multiple views into distributions and computes a single contrastive loss between them.
@@
To address this, we propose a novel similarity function that models the joint structure among all augmented views of data. We estimate a probability distribution on the unit hypersphere and compute similarity via negative divergence between distributions, see Figure 1. We use the von Mises–Fisher (vMF) distribution, a widely used Gaussian-like distribution on the unit hypersphere (Mardia and Jupp 2009; Fisher, Lewis, and Embleton 1993; Kitagawa and Rowley 2022; Scott, Gallagher, and Mozer 2021; Du et al. 2024). vMF distribution is parameterized by mean direction $\mu$ and concentration $\kappa$, which respectively represent the central tendency and joint structure of features. Notably, in the two-view case, our method recovers cosine similarity in the InfoNCE loss (Oord, Li, and Vinyals 2018), linking pairwise and multi-view contrastive frameworks.
@@
  • Establishing the relation to the original InfoNCE loss. We show that, when restricted to two views under specific condition, our similarity recovers cosine similarity in InfoNCE loss, thus offering the unified view of pairwise and multi-view contrastive methods.
  • Temperature-free. We show the dependency of the temperature term in cosine similarity and show that, for DSF, temperature tuning is unnecessary.
  • Superior Experimental Results. We validate the effectiveness of DSF through extensive experiments on kNN classification, linear evaluation, transfer learning, and distribution shift, as well as resource analysis in terms of GPU memory and training time, consistently outperforming both baseline and existing multi-view contrastive learning methods.
The remainder of this paper is structured as follows. Section 2 reviews related work on contrastive frameworks with both two-view ($M=2$) and multi-view ($M\geq 3$) settings. Section 3 provides an overview of CL and its multi-view extensions. Section 4 presents our proposed formulation. Section 5 describes the experimental setup and results. Finally, Section 6 concludes the paper and discusses potential directions for future research.
2 Related Works
@@
[…]
2.1 Original Contrastive Learning
CL appears in many self-supervised learning methods. SimCLR (Chen et al. 2020a) leverages the InfoNCE loss (Oord, Li, and Vinyals 2018) to pull together representations of two augmented views while pushing apart all other samples. MoCo (He et al. 2020) introduces a momentum encoder and a dynamic dictionary (negative queue) to maintain a large and consistent set of negative samples. DCL (Yeh et al. 2022) decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable and high performance training with smaller batches and without a momentum encoder. MoCo v3 (Chen, Xie, and He 2021) explores CL using a transformer-based model, specifically the Vision Transformer (ViT) (Dosovitskiy et al. 2020).
2.2 Contrastive Learning with $M\geq 3$ views
Existing $M\geq 3$ multi‐view CL methods have broadly fallen into two categories. The first category shapes the loss by averaging pairwise terms across views. (Shidani et al. 2024) generates $M$ views per a sample and aggregates their pairwise losses either arithmetically or geometrically. The second category is a method that deals with view representations at the feature level before computing the final objective. FastSiam (Pototzky, Sultan, and Schmidt-Thieme 2022) averages the other (M1M-1) view features to construct a target of each view, and M3G (Piran et al. 2024) adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost. In contrast to these approaches, DSF integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
3 Preliminaries
@@ 3.1 Contrastive Learning
CL has emerged as a powerful paradigm in unsupervised and self-supervised learning, aiming to learn representations by distinguishing positive and negative pairs. Positive pair refers to semantically similar pair whereas negative pair to dissimilar ones. The core idea is to map positive samples closer while pushing negative samples apart in the embedding space. Given a positive pair and a set of negative samples, InfoNCE loss encourages the model to assign high similarity to the positive pair relative to the negatives. The loss can be formulated as follows:
$L^{\mathrm{cos}}_{i,\text{InfoNCE}}=-\log\frac{e^{(\mathrm{sim}(z_{i},z^{+}_{i})/\tau)}}{e^{(\mathrm{sim}(z_{i},z^{+}_{i})/\tau)}+\sum\limits_{j\neq i}^{K}e^{(\mathrm{sim}(z_{i},z^{-}_{j})/\tau)}},$ (1)
where $z^{+}_{i}$ is the embedding of the positive pair respect to $z_{i}$ and $z^{-}_{j}$s are embeddings of K1K-1 negative pairs. $\mathrm{sim}(\cdot,\cdot)$ is a similarity measure (e.g., dot product) and $\tau$ is a temperature parameter.
@@
vMF distribution is defined on the (p1)(p\!-\!1)-dimensional unit hypersphere $\mathbb{S}_{p-1}$ and models the probability density of unit vectors with a specified mean direction $\mu$ and concentration parameter $\kappa$. Its probability density function for a unit vector $x\in\mathbb{S}_{p-1}$ is:
$f_{p}(x;\mu,\kappa)=C_{p}(\kappa)\exp(\kappa\mu^{T}x),$ (2)
where $C_{p}(\kappa)$ is the normalization constant. Higher $\kappa$ values yield distributions more concentrated around $\mu$, while $\kappa=0$ corresponds to a uniform distribution on the unit hypersphere.
3.3 Notation
We denote the batch size, number of views, and feature dimension by $B$, $M$, and $p$, respectively. Each encoder is given $m$ views, where $m$ indicates the number of augmented views per encoder, resulting in a total of $M=2m$ views.
Let $x$ be an image data, and $x^{l}$ is denoted by its $l$-th augmented view. Given an encoder $f_{\theta}(\cdot)$, the corresponding feature on the unit hypersphere is computed as $z^{l}=f_{\theta}(x^{l})/\|f_{\theta}(x^{l})\|$, $l=1,\cdots,m$. In CL, there are $2\cdot B$ features $z_{i}$, $i=1,\cdots,2B$. $z_{i}$ and $z_{B+i}$ are positively correlated. In multi-view CL, there are $2m\cdot B$ features $z_{i}^{l}$, $i=1,\cdots,B$ and $l=1,\cdots,m$. $\{z_{i}^{l}\}_{l=1}^{m}$ and $\{z_{B+i}^{l}\}_{l=1}^{m}$ are positively correlated.
The Kullback–Leibler (KL) divergence is written as $D_{\mathrm{KL}}(\cdot\|\cdot)$, and we denote a vMF distribution by the calligraphic symbol $\mathcal{D}$.
@@ 4 Divergence-based similarity function
In this section, we describe DSF in detail and demonstrate its relationship to cosine similarity, as well as its inherent temperature-free property.
4.1 Estimation of vMF Distribution
First, we describe how to construct vMF distributions from multiple augmented views of an image.
Given data $x$, we sample $M$ augmented views and divide them into two groups of $m$ samples each. These are passed through the encoder to obtain unit-normalized feature vectors. Each group is then used to estimate a vMF distribution, yielding two vMFs per a data.
Let $z^{1},\dots,z^{m}$ be the feature vectors from one group. The mean direction $\mu$ and concentration $\kappa$ are estimated as:
@@ 4.1 Estimation of vMF Distribution
As $A_{p}^{-1}$ is intractable, we use the approximation from (Banerjee et al. 2005):
$\kappa=\frac{\bar{R}(p-\bar{R}^{2})}{1-\bar{R}^{2}}.$ (5)
Let $g_{\text{est}}$ be the estimation function mapping features to vMF parameters:
$g_{\text{est}}(z^{1},\dots,z^{m})=(\mu,\kappa),$
and define the resulting vMF distribution as:
$\mathcal{D}\left(g_{\text{est}}(z^{1},\dots,z^{m})\right)=\mathcal{D}(\mu,\kappa).$ (6)
For data $i$, we denote:
$\mathcal{D}\left(g_{\text{est}}(z_{i}^{1},\dots,z_{i}^{m})\right)=\mathcal{D}(\mu_{i},\kappa_{i})=\mathcal{D}_{i}.$ (7)
@@ 4.2 Stabilization of the Concentration Parameter $\kappa$
In this section, we describe a stabilization strategy for the concentration parameter $\kappa$, which is critical for preventing numerical issues during training.
As training progresses, $\kappa$ may grow excessively due to:
  • High feature dimensionality: Larger feature dimensions $p$ result in sharper increase in $\kappa$, especially with high-dimensional embeddings (e.g., $D=128$). The problem is further exacerbated when larger models are used.
  • Feature concentration: As representations improve, features become more aligned and $\bar{R}$ approaches 1, further sharp increasing $\kappa$.
Though theoretically valid, such growth causes numerical instability in divergence computation. To mitigate this, we adopt two strategies:
@@
    Dimension-based normalization: Normalize $\kappa$ by the feature dimension $p$, since $\kappa$ grows roughly linearly with respect to $p$ when $\bar{R}$ is fixed (see Eq. 5).
@@
    Scaling mean resultant length: Prevent $\bar{R}$ from approaching 1 by applying a scaling factor $\lambda_{\bar{R}}=0.95$:
@@
These techniques are applied after estimating the vMF parameters, and they help stabilize training and prevent divergence-related issues in high-dimensional settings. Further details are provided in Appendix C.3, Figure 6.
Dataset Base Model Method Epoch kNN (↑) Linear Evaluation (↑) mCE* (↓)
ImageNet-100 MoCo v3 - 300 74.08 83.42 36.67
MoCo v3 Loss_avg 100 76.62 84.06 38.27
MoCo v3 Feature_avg 100 76.94 83.82 39.42
MoCo v3 DSF 100 79.00 85.80 33.15
CIFAR-10 MoCo - 800 88.44 88.10 19.03
MoCo Loss_avg 250 88.28 88.58 19.90
MoCo Featrue_avg 250 87.47 88.02 19.80
MoCo DSF 250 90.04 91.21 16.20
Table 1: Comparison of kNN, Linear Evaluation and mCE Results Under Equivalent Resource Settings. All experiments are conducted under the same GPU memory and time constraints. The number of views is fixed at $M=2m=8$. The mean Corruption Error(mCE) was evaluated on CIFAR-10-C and ImageNet-100-C, with all models trained solely on clean(uncorrupted) data.